A double tennis court is a rectangle 36 feet wide and 78 feet long. If two players are standing in diagonally opposite corners, about how far apart are they, to the nearest tenth of a foot.
Answer by addingup(3677)

The line cutting the rectangle diagonally gives you two identical triangles and the dividing line is the hypotenuse. So we apply the Pythagorean formula:
a^2+b^2 = c^2
36^2+78^2 = c^2
1296+6084 = 7380 take the square root: sqrt7380 = 85.9
The players are standing about 85.9 feet apart